Quadratic Equations
20) The perimeter of a rectangle is 8m and its area is 2m^2. Find the length and width of the rectangle to the nearest tenth of a metre.
Thank you

Answer by edjones(8007)

A) L*w=2
.
2L+2w=8
L+w=4 divide equation by 2.
L=4-w
.
A) w(4-w)=2
4w-w^2=2
w^2-4w+2=0
w^2-4w =-2
w^2-4w+4=-2+4 complete the square
(w-2)^2=2
w-2=+-sqrt(2)
w=2+-sqrt(2)
w=2+sqrt(2), L=2-sqrt(2) or vice versa
.
Ed

.